OpenStudy (erich):

Solve for X px^2-qx+2=0

OpenStudy (erich):

so then I did x= q + or - the squareroot of -q^2 - 4 (p) (2) all over 2p

OpenStudy (erich):

is this as simplified as it can go?

OpenStudy (anonymous):

\[\frac{q \pm \sqrt{q^2-8p}}{2p}\]So, yeah. But your -q^2 should be (-q)^2=q^2